OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (jdoe0001):

\(\bf \textit{angle between two vectors }\\ \quad \\ cos(\theta)=\cfrac{u \cdot v}{||u||\ ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\\ \quad \\ \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right)\)

OpenStudy (anonymous):

I did that and got 0.99

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

i'll show you my work if you want

OpenStudy (jdoe0001):

hmm what did you get for the dot product?

OpenStudy (anonymous):

(-5)(-4) + (-4)(-3) = 20+ 12 = 32

OpenStudy (jdoe0001):

ok.... lemme check

OpenStudy (anonymous):

ohhh i think i know what i did

OpenStudy (anonymous):

one sec im going to redo it

OpenStudy (anonymous):

is it 1.8?

OpenStudy (jdoe0001):

yeap 1.789910 degrees, thus \(\large 1.8^o\)

OpenStudy (anonymous):

thanks so much!!! :)

OpenStudy (jdoe0001):

yw